# On Efficient Variants of Segment Anything Model: A Survey

### v2

@@ Abstract
The Segment Anything Model (SAM) is a foundational model for image segmentation tasks, known for its strong generalization across diverse applications. However, its impressive performance comes with significant computational and resource demands, making it challenging to deploy in resource-limited environments such as edge devices. To address this, a variety of SAM variants have been proposed to enhance efficiency while keeping accuracy. This survey provides the first comprehensive review of these efficient SAM variants. We begin by exploring the motivations driving this research. We then present core techniques used in SAM and model acceleration. This is followed by a detailed exploration of SAM acceleration strategies, categorized by approach, and a discussion of several future research directions. Finally, we offer a unified and extensive evaluation of these methods across various hardware, assessing their efficiency and accuracy on representative benchmarks, and providing a clear comparison of their overall performance.
@@ 1 Introduction
Recently, a novel foundation model for general image segmentation, the Segment Anything Model (SAM), was proposed by Meta [62]. Being fully trained on their proposed SA-1B dataset, which consists of more than one billion masks and eleven million images, with the task of achieving valid segmentation given any prompt (e.g. points, boxes, masks and text), SAM is able to well generalize to a wide range of downstream tasks (e.g. edge detection, object proposal and instance segmentation) with proper prompts. Shortly after its emergence, SAM has garnered significant attention from the research community and results in a surge of related works exploring SAM’s generalization ability in various scenarios [88, 16, 17], including different image segmentation tasks [107, 64, 2, 142], video analysis tasks [166, 149, 86, 48] and 3D vision tasks [112, 30, 146]. With the huge success of SAM, the upgraded Segment Anything Model 2 (SAM 2) [105] is further proposed, aiming for efficient segmentation in both images and videos. SAM 2 introduces the streaming memory mechanism to extend SAM’s ability to videos. It is trained on both the SA-1B dataset and SA-V dataset which is their newly collected video segmentation dataset. As a result, SAM 2 naturally holds the powerful generalization and surpasses SAM when handling segmentation tasks [111, 147, 152, 125, 80].
@@ 2.1.3 Application
One of the major applications of SAM is medical image segmentation. According to [169], works in this area can be classified into two types. Some aim at testing SAM’s segmentation performance in CT images [107, 53], MRI images [93], pathological images [27], and so on. Others focus on improving the adaptation of SAM to these tasks by fine-tuning [88, 71], auto-prompting [55, 110] or framework modification [168]. Furthermore, works like [64, 39, 90] attempt to improve medical SAM methods’ efficiency. SAM is also applied to object detection across a diverse range of real-world scenarios [161, 58], including crack detection [2] and crater detection [40] in civil infrastructure defect assessment, crop disease and pest detection [70] in agriculture, anomaly detection [11] and remote sensing [96, 131]. Moreover, Segment Anything has been adapted to Edit Everything [142], Inpaint Anything [153], and Caption Anything [133] to handle image editing tasks.
@@
Apart from image segmentation tasks, SAM has been widely applied to various video tasks [163]. A large portion of the work concentrates on two basic tasks: video object segmentation (VOS) [166, 170, 21] and video object tracking (VOT) [149, 22, 104]. Researchers also explore SAM’s application in tasks related to generation, for example, video super-resolution [86] and video dataset annotation generation [48, 26]. Besides these, SAM is further utilized as an intermediate tool in video editing tasks [175, 138]. Beyond 2D tasks, SAM is also extended to the 3D vision field. Shen et al. [112] applies SAM to 3D reconstruction and Dong et al. [30] applies it to 3D point cloud segmentation. The recent work [146] aims at achieving real-time segmentation for any 3D thing in an online setting. For the recently proposed SAM 2, some studies have already explored its application in both image and video tasks. A popular trend is to apply SAM 2 to medical images and video tasks [167]. Works like [111, 29, 147] evaluate SAM 2’s performance in medical images in both 2D and 3D modalities, while some others [152, 85] test it on surgical video segmentation tasks. Researchers are also seeking strategies to adapt SAM 2 to medical tasks better [148, 178]. Besides these, SAM 2 has been applied to some specific image segmentation tasks like digital pathology semantic segmentation [165], mesh part segmentation [121], and solar panels segmentation [103]. Moreover, several works [125, 80] have utilized SAM 2 on the challenging Large-scale Video Object Segmentation (LSVOS) task and achieved good results.
@@ 2.1.4 Limitation
Although SAM has demonstrated promising performance across various tasks, it still faces two key challenges in practical applications: 1) it often struggles to predict complete masks for fine structures, leading to imprecise boundaries; and 2) it is not real-time and remains resource-intensive, particularly when using heavy image encoders like ViT-H. To address these issues, works such as [61, 56] aims to improve mask quality by utilizing high-resolution images, while others [173, 159, 176, 143] focus on creating more efficient architectures to reduce SAM’s time and resource consumption. Previous surveys [88, 170, 112] have explored recent advances in enhancing SAM for higher-quality results. In this survey, we focus specifically on efforts to improve SAM’s efficiency.
2.2 Efficient Backbone
[…]
Table 1: Parameters of SAM’s encoders.
@@ 2.3.3 Pruning
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=16em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, where level=4text width=18em,font=,, where level=5text width=12em,font=,, [Efficient SAM variants, ver [Accelerating SegAny3.1), fill=green!10 [Training from Scratch3.1.1), fill=blue!10 [SAM-different architectures, fill=yellow!10 [FastSAM [173], fill=gray!10] [SqueezeSAM [126], fill=gray!10] ] [SAM-like architectures, fill=yellow!10 [EfficientSAM [143], fill=gray!10] [RAP-SAM [145], fill=gray!10] [SAM 2 [105], fill=gray!10] ] ] [Knowledge Distillation3.1.2), fill=blue!10 [With Light-weight ViT Encoders, fill=yellow!10 [MobileSAM [159], fill=gray!10] [ESAM [174], fill=gray!10] [TinySAM [115], fill=gray!10] ] [With Pure CNN Encoders, fill=yellow!10 [NanoSAM [95], fill=gray!10] [RepViT-SAM [130], fill=gray!10] [EdgeSAM [176], fill=gray!10] ] [With Attention-modified Encoders, fill=yellow!10 [EfficientViT-SAM [171], fill=gray!10] [FastSAM3D [113], fill=gray!10] [SAM-Lightening [116], fill=gray!10] [RWKV-SAM [99], fill=gray!10] ] ] [Quantization3.1.3), fill=blue!10 [TinySAM [115], fill=gray!10] [PTQ4SAM [87], fill=gray!10] ] [Pruning3.1.4), fill=blue!10 [SlimSAM [20], fill=gray!10] ] [Code Refactorization3.1.5), fill=blue!10 [SAMFast [100], fill=gray!10] ] ] [Accelerating SegEvery3.2), fill=green!10 [Panoptic Segmentation Strategy, fill=blue!10 [FastSAM [173], fill=gray!10, tier=number] ] [Efficient Sampling Strategy, fill=blue!10 [MobileSAMV2 [160], fill=gray!10, tier=number] [TinySAM [115], fill=gray!10] [LiteSAM [36], fill=gray!10] ] ] ]
@@ 3 Efficient Variants of SAM
This section reviews the efforts to develop lightweight, efficient SAM-like models that have emerged since SAM gained prominence. These works aim to reduce the model’s high computational cost and enable efficient performance, while preserving SAM’s robust segmentation capabilities and generalization strength. As outlined in Section 2.1.2, SAM addresses two primary tasks including Segment Anything (SegAny) and Segment Everything (SegEvery). Accordingly, we discuss the research aimed at improving each task separately: Section 3.1 focuses on accelerating the SegAny task, and Section 3.2 covers efforts to accelerate the SegEvery task. Notably, some methods are applicable to both tasks, and we discuss these contributions individually. Additionally, we categorize all models into different classes based on the techniques they employ, and the taxonomy of methodologies is presented in Fig. 3. Finally, in Section 3.3, we outline the potential directions for future research in this area.
@@ 3.1.1 Training from Scratch
Building on the successful application of CNNs in FastSAM, Varadarajan et al. introduced SqueezeSAM [126], which further replaces SAM’s Transformer-based architecture with a U-Net structure [106]. U-Net is composed of an encoder for feature extraction and a decoder for information recovery. SqueezeSAM retains the general U-Net architecture but incorporates two Transformer layers at the lowest scale of the U-Net to strike a balance between speed and accuracy. Additionally, SqueezeSAM features several micro-level optimizations, such as capping the output channels at 256, using BatchNorm [57] instead of LayerNorm [5] for efficiency, and introducing skip connections [49] between the encoder and decoder. A unique challenge for SqueezeSAM lies in handling prompts. Unlike SAM, where prompt tokens are utilized at the decoding stage, SqueezeSAM adopts an early fusion strategy, adding encoded prompts as additional input channels before feeding them into the encoder. The model is trained from scratch using the SA-1B dataset, with data augmentation techniques addressing low-quality data issues. SqueezeSAM is primarily designed for deployment in photography applications, where efficient interactive segmentation is needed. As depicted in Fig. 5, its workflow involves generating an initial mask of the salient object, followed by fine-grained segmentation refined by user clicks.
@@
Instead of introducing an entirely new network, EfficientSAM [143] retains SAM’s original architecture but replaces the image encoder. They use ViT-tiny or ViT-small as a lightweight encoder and re-train them from scratch, leveraging the SAM-based Masked Image (SAMI) pretraining strategy. SAMI is adapted from the Masked AutoEncoder (MAE) framework [50], which was initially used to pretrain SAM’s original image encoder. SAMI follows an Encoder-Decoder pipeline: the encoder generates latent feature representations from unmasked patches, while the decoder reconstructs the missing embeddings of masked patches. This process is supervised by reconstruction loss, comparing the embeddings produced by SAM’s ViT-H encoder with those generated by the SAMI pipeline. After pretraining, the lightweight encoder is extracted from the SAMI pipeline and integrated with the rest of SAM’s components, forming EfficientSAM. The final step involves fine-tuning the entire model on the SA-1B dataset for further alignment and refinement. SAMI is a general pretraining approach that can be applied to train any backbone for SAM variants. The overall structure of SAMI and EfficientSAM is illustrated in Fig. 6.
@@
Xu et al. propose the RAP-SAM [145], a model designed to achieve real-time, all-purpose segmentation, including panoptic segmentation (PS), video instance segmentation (VIS), and interactive segmentation (equivalent to the SegAny task). RAP-SAM retains SAM’s basic Encoder-Decoder architecture but incorporates more efficient components to enhance performance. For the encoder, RAP-SAM combines Feature Pyramid Networks (FPN) [75] with deformable convolutions [24] to extract features from both images and videos, while using a prompt encoder to embed visual prompts. In the decoder, RAP-SAM employs a three-stage pipeline utilizing novel pooling-based dynamic convolutions to refine the mask tokens. The tokens generated in each stage, along with the feature maps from the encoder, serve as inputs. These inputs are first processed by dynamic convolutions [19], followed by refinement using Multi-Head Self Attention (MHSA) and a Feed Forward Network (FFN). After the decoder, two additional prompt adapters are introduced to enhance interaction between the visual prompts and segmentation tokens. The final masks are generated by computing the inner product between the updated tokens and updated prompts. The architecture of RAP-SAM is illustrated in Fig. 7.
Figure 7: The architecture of RAP-SAM[145]. The encoder embeds images and videos to feature maps and encodes visual prompts as queries (tokens). The multi-stage decoder is to refine queries by interacting with feature maps. In the end, the feature maps and refined queries are fed into adapters to generate masks and classes.
Recently, Ravi et al. introduced the Segment Anything Model 2 (SAM 2) [105], an extension of the original SAM. The objective of SAM 2 is to deliver high-quality, real-time promptable segmentation across both images and videos. In image segmentation tasks, SAM 2 is reported to achieve higher accuracy and a 6× improvement in efficiency compared to the original SAM. This significant advancement is largely attributed to its efficient image encoder, Hiera [108], a hierarchical ViT that has been simplified from MViTv2 [67] by removing redundant components and utilizing the MAE framework for training. Hiera is a streamlined, purely transformer-based architecture that runs faster and delivers better accuracy than traditional ViTs in both image and video tasks.

### v3

@@ Abstract
The Segment Anything Model (SAM) is a foundational model for image segmentation tasks, known for its strong generalization across diverse applications. However, its impressive performance comes with significant computational and resource demands, making it challenging to deploy in resource-limited environments such as edge devices. To address this, a variety of SAM variants have been proposed to enhance efficiency while keeping accuracy. This survey provides the first comprehensive review of these efficient SAM variants. We begin by exploring the motivations driving this research. We then present core techniques used in SAM and model acceleration. This is followed by a detailed exploration of SAM acceleration strategies, categorized by approach, and a discussion of several future research directions. Finally, we offer a unified and extensive evaluation of these methods across various hardware, assessing their efficiency and accuracy on representative benchmarks, and providing a clear comparison of their overall performance. To complement this survey, we summarize the papers and codes related to efficient SAM variants at https://github.com/bhllx/On-Efficient-Variants-of-Segment-Anything-Model.
keywords:
Segment Anything Model, Efficient Backbone, Light-Weight Architecture, Model Compression, Model Acceleration.
@@ 1 Introduction
Recently, a novel foundation model for general image segmentation, the Segment Anything Model (SAM), was proposed by Meta [19]. Being fully trained on their proposed SA-1B dataset, which consists of more than one billion masks and eleven million images, with the task of achieving valid segmentation given any prompt (e.g. points, boxes, masks and text), SAM is able to well generalize to a wide range of downstream tasks (e.g. edge detection, object proposal and instance segmentation) with proper prompts. Shortly after its emergence, SAM has garnered significant attention from the research community and results in a surge of related works exploring SAM’s generalization ability in various scenarios [20, 21, 22], including different image segmentation tasks [23, 24, 25, 26], video analysis tasks [27, 28, 29, 30] and 3D vision tasks [31, 32, 33]. With the huge success of SAM, the upgraded Segment Anything Model 2 (SAM 2) [34] is further proposed, aiming for efficient segmentation in both images and videos. SAM 2 introduces the streaming memory mechanism to extend SAM’s ability to videos. It is trained on both the SA-1B dataset and SA-V dataset which is their newly collected video segmentation dataset. As a result, SAM 2 naturally holds the powerful generalization and surpasses SAM when handling segmentation tasks [35, 36, 37, 38, 39, 40].
@@
Table 1: Efficient SAM Variants: Backbone and Key Features.
Variant Backbone Key Features
FastSAM YOLOv8-Seg (CNN) All-instance segmentation then prompt-guided selection.
SqueezeSAM CNN U-Net based encoder-decoder architecture.
EfficientSAM ViT-Tiny/Small Retrain lightweight ViT encoder with SAM leveraged.
RMP-SAM Light CNNs Unified dynamic decoder with multi-task adapters.
MobileSAM TinyViT (ViT) Encoder-only distillation from SAM.
Light HQ-SAM TinyViT (ViT) Merge MobileSAM’s efficiency with HQ-SAM’s quality.
ESAM EffiFormerV2 (ViT) Holistic distillation from an expert model.
NanoSAM ResNet (CNN) Knowledge distillation and TensorRT acceleration.
RepViT-SAM RepViT (CNN) ViT-inspired CNN design and efficient distillation.
EdgeSAM RepViT (CNN) Prompt-in-loop distillation and embedding of priors.
EfficientViT-SAM EffiViT (ViT) Apply linear ReLU attention to cut computation.
FastSAM3D ViT-Tiny Extends SAM to 3D with sparse flash attention.
SAM-Lightening ViT-Tiny Dynamic layer-wise distillation for faster inference.
RWKV-SAM MBConv + VRWKV Integrate linear RWKV operation for efficiency.
TinySAM TinyViT (ViT) Hierarchical sampling with full-stage distillation.
PTQ4SAM ViT Post-training quantization with adaptive granularity.
PQ-SAM ViT Group activations for efficient quantization.
SlimSAM ViT Structured pruning using disturbed Taylor importance.
SuperSAM ViT One-shot NAS combined with pruning.
SAMfast ViT Native PyTorch optimizations yield an 8× speedup.
Lite-SAM Lite-ViT (Hybrid) Multi-scale pooling and AutoPPN for prompt generation.
MobileSAMv2 ViT Object-aware prompt sampling.
AoP-SAM ViT Adaptive prompt prediction with coarse-to-fine sampling.
@@ 2.1.3 Application
One of the major applications of SAM is medical image segmentation. According to [52], works in this area can be classified into two types. Some aim at testing SAM’s segmentation performance in CT images [23, 60], MRI images [61], pathological images [62], and so on. Others focus on improving the adaptation of SAM to these tasks by fine-tuning [20, 63], auto-prompting [64, 65, 66] or framework modification [67]. Furthermore, as these SAM-based works also suffered from high latency and heavy computational cost, improving medical SAM methods’ efficiency has increasingly gained attention in the community [68]. Several lightweight architectures [24, 69, 70, 71, 72] have emerged, achieving both high-speed and high-quality segmentation across various medical imaging modalities, most of which distilled MedSAM’s [20] image encoder into a lighter one to improve efficiency while maintaining the accuracy by subsequent fine-tuning. Optimization techniques during inference, for instance, exporting models to OpenVINO format and implementing an embedding caching mechanism [24, 69], could further speed up runtime efficiency, enabling variants to deploy on resource-constrained environments even without GPU.
@@
SAM is also applied to object detection across a diverse range of real-world scenarios [50, 73], including crack detection [25] and crater detection [74] in civil infrastructure defect assessment, crop disease and pest detection [75] in agriculture, anomaly detection [76] and remote sensing [77, 78, 79]. Moreover, Segment Anything has been adapted to Edit Everything [26], Inpaint Anything [80], and Caption Anything [81] to handle image editing tasks. Additionally, SAM is applied to provide priors for high-quality multi-modality image fusion [82]. Some researchers focus on exploring the uncertainty in SAM [83, 84]. For instance, UncertainSAM [83] has proposed effective uncertainty quantization for SAM, and UNICL-SAM[84] further leverage the uncertainty to guide segmentation.
@@
Apart from image segmentation tasks, SAM has been widely applied to various video tasks [53]. A large portion of the work concentrates on two basic tasks: video object segmentation (VOS) [27, 85, 86, 87] and video object tracking (VOT) [28, 88, 89]. Researchers also explore SAM’s application in tasks related to generation, for example, video super-resolution [29] and video dataset annotation generation [30, 90]. Besides these, SAM is further utilized as an intermediate tool in video editing tasks [91, 92]. Beyond 2D tasks, SAM is also extended to the 3D vision field. Shen et al. [31] applies SAM to 3D reconstruction and Dong et al. [32] applies it to 3D point cloud segmentation. The recent work [33] aims at achieving real-time segmentation for any 3D thing in an online setting. For the recently proposed SAM 2, some studies have already explored its application in both image and video tasks. A popular trend is to apply SAM 2 to medical images and video tasks [54]. Works like [35, 93, 36] evaluate SAM 2’s performance in medical images in both 2D and 3D modalities, while some others [38, 94] test it on surgical video segmentation tasks. Researchers are also seeking strategies to adapt SAM 2 to medical tasks better [95, 96]. Besides these, SAM 2 has been applied to some specific image segmentation tasks like digital pathology semantic segmentation [97], mesh part segmentation [98], solar panels segmentation [99], and 3D instance segmentation [100]. Moreover, several works [101, 102, 103, 39, 40] have improved SAM 2’s performance on the VOT task, the Referring Video Object Segmentation (RVOS) task and the challenging Large-scale Video Object Segmentation (LSVOS) task.
@@ 2.1.4 Limitation
Although SAM has demonstrated promising performance across various tasks, it still faces two key challenges in practical applications: 1) it often struggles to predict complete masks for fine structures, leading to imprecise boundaries; and 2) it is not real-time and remains resource-intensive, particularly when using heavy image encoders like ViT-H. To address these issues, some works aim to improve mask quality by utilizing high-resolution images [104, 105], integrating images and prompts [106], or introducing optimized prompts [107, 108], while others [41, 43, 46, 109] focus on creating more efficient architectures to reduce SAM’s time and resource consumption. Some recent works, such as  [110], have also started to find a better balance between model’s accuracy and efficiency. Previous surveys [20, 85, 31] have explored recent advances in enhancing SAM for higher-quality results. In this survey, we focus specifically on efforts to improve SAM’s efficiency.
Table 2: Parameters of SAM’s encoders. Results in this table refer to [43].
@@ 2.3.1 Knowledge Distillation
Discussion: Compared to training model fully from scratch, knowledge distillation usually requires fewer training resources, making it a relatively economical choice to build up lightweight and effective variants of large models, and thus being widely leveraged in accelerating SAM. Moreover, with further fine-tuning on task-specific and smaller datasets, the distilled model can be effectively adapted to several downstream tasks, even in the resource-constrained environment which is not suited for deploying the heavy powerful model. However, distilling knowledge introduces additional supervision (i.e., the student model learns from both the soft labels of the teacher model and the ground truth labels of the task), making the training process more complex, which may lead to optimization difficulty and additional cost. As a result, works accelerating SAM based on KD have all carefully designed the distilling pipeline, with expectation to improve the quality of distillation.
@@ 2.3.2 Quantization
Discussion: Post-Training Quantization is much more broadly investigated in lightening SAM, as it allows for directly applying quantization to the pre-trained foundation model with a few data for calibration, which is flexible and efficient, while on the contrary, Quantization-Aware Training requires re-training the model on entire training set, with heavier training cost. However, PTQ-based light models usually suffer a bigger drop in accuracy compared to QAT-based ones, especially for those not explicitly trained to be quantization-aware. In the practice of applying quantization on SAM while maintaining a reasonable level of performance, the extremely complex and asymmetric distribution rooted in SAM has made the process much more challenging.
@@ 2.3.3 Pruning
Discussion: Weights, channels and even layers in a large model are not equally important to contribute to the accuracy, making model pruning an intuitive yet effective method to reduce the inner redundancy. Nevertheless, the pruned model may not always achieve the expected speedup on certain hardware. Another potential problem is that pruning aggressively can sometimes lead to the loss of generalization, which is unacceptable when lightening foundation models like SAM, leading to the demand for criteria that comprehensively evaluates weights’ significance.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=16em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, where level=4text width=18em,font=,, where level=5text width=12em,font=,, [Efficient SAM variants, ver [Accelerating SegAny3.1), fill=green!10 [Training from Scratch3.1.1), fill=blue!10 [SAM-different architectures, fill=yellow!10 [FastSAM [41], fill=gray!10] [SqueezeSAM [42], fill=gray!10] ] [SAM-like architectures, fill=yellow!10 [EfficientSAM [109], fill=gray!10] [RMP-SAM [150], fill=gray!10] [SAM 2 [34], fill=gray!10] ] ] [Knowledge Distillation3.1.2), fill=blue!10 [With Light-weight ViT Encoders, fill=yellow!10 [MobileSAM [43], fill=gray!10] [Light HQ-SAM [104], fill=gray!10] [ESAM [121], fill=gray!10] [TinySAM [44], fill=gray!10] ] [With Pure CNN Encoders, fill=yellow!10 [NanoSAM [138], fill=gray!10] [PicoSAM2 [151], fill=gray!10] [RepViT-SAM [45], fill=gray!10] [EdgeSAM [46], fill=gray!10] ] [With Attention-modified Encoders, fill=yellow!10 [EfficientViT-SAM [122], fill=gray!10] [FastSAM3D [128], fill=gray!10] [SAM-Lightening [129], fill=gray!10] [RWKV-SAM [117], fill=gray!10] ] ] [Quantization3.1.3), fill=blue!10 [TinySAM [44], fill=gray!10] [PTQ4SAM [47], fill=gray!10] [PQ-SAM [152], fill=gray!10] ] [Pruning3.1.4), fill=blue!10 [SlimSAM [48], fill=gray!10] [SuperSAM [153], fill=gray!10] ] [Code Refactorization3.1.5), fill=blue!10 [SAMFast [154], fill=gray!10] ] ] [Accelerating SegEvery3.2), fill=green!10 [Panoptic Segmentation Strategy, fill=blue!10 [FastSAM [41], fill=gray!10, tier=number] ] [Efficient Sampling Strategy, fill=blue!10 [MobileSAMV2 [155], fill=gray!10, tier=number] [TinySAM [44], fill=gray!10] [LiteSAM [124], fill=gray!10] [AoP-SAM [156], fill=gray!10] ] ] ]
@@ 2.3.4 Low-Rank Factorization
Discussion: Sparse matrices derived from low-rank factorization, will take less memory usage and computational cost during forward pass, slimming and accelerating the vanilla model. However, the implementation of some decomposition operations, for instance the SVD, yet being stable and effective in fully connected layers [164], can be computationally expensive especially for large matrices [165], which may hinders its application in SAM. Moreover, determining the best matrix decomposition rank remains to be the crucial challenge for applying factorization approaches. Furthermore, similar to pruning-based methods, decomposing matrices into low rank may result in loss of detailed information, undermining model’s generalization performance. Even though, some recent works applying low-rank factorization in LLMs [166, 167, 168], exploiting the decomposition according to activation distribution or in specific Transformer layers, are proved to be effective, which may provide inspiration to researchers to further implement it on SAM.
@@ 3 Efficient Variants of SAM
This section reviews the efforts to develop lightweight, efficient SAM-like models that have emerged since SAM gained prominence. These works aim to reduce the model’s high computational cost and enable efficient performance, while preserving SAM’s robust segmentation capabilities and generalization strength. As outlined in Section 2.1.2, SAM addresses two primary tasks including Segment Anything (SegAny) and Segment Everything (SegEvery). Accordingly, we discuss the research aimed at improving each task separately: Section 3.1 focuses on accelerating the SegAny task, and Section 3.2 covers efforts to accelerate the SegEvery task. Notably, some methods are applicable to both tasks, and we discuss these contributions individually. To help readers have a quick view of the efficient SAM variants, we provide Tab. 1, summarizing their backbone choices and key features. Additionally, we categorize all models into different classes based on the techniques they employ, and the taxonomy of methodologies is presented in Fig. 3. Finally, in Section 3.3, we outline the potential directions for future research in this area.
@@ 3.1.1 Training from Scratch
Building on the successful application of CNNs in SAM, Varadarajan et al. introduced SqueezeSAM [42], which further replaces SAM’s Transformer-based architecture with a U-Net structure [111]. U-Net is composed of an encoder for feature extraction and a decoder for information recovery. SqueezeSAM retains the general U-Net architecture but incorporates two Transformer layers at the lowest scale of the U-Net to strike a balance between speed and accuracy. Additionally, SqueezeSAM features several micro-level optimizations, such as capping the output channels at 256, using BatchNorm [171] instead of LayerNorm [172] for efficiency, and introducing skip connections [112] between the encoder and decoder. A unique challenge for SqueezeSAM lies in handling prompts. Unlike SAM, where prompt tokens are utilized at the decoding stage, SqueezeSAM adopts an early fusion strategy, adding encoded prompts as additional input channels before feeding them into the encoder. The model is trained from scratch using the SA-1B dataset, with data augmentation techniques addressing low-quality data issues. SqueezeSAM is primarily designed for deployment in photography applications, where efficient interactive segmentation is needed. As depicted in Fig. 5, its workflow involves generating an initial mask of the salient object, followed by fine-grained segmentation refined by user clicks.
@@
Different from FastSAM and SqueezeSAM, which introduce entirely new structures, EfficientSAM [109] retains SAM’s original architecture but replaces the image encoder. They use ViT-tiny or ViT-small as a lightweight encoder and re-train them from scratch, leveraging the SAM-based Masked Image (SAMI) pretraining strategy. SAMI is adapted from the Masked AutoEncoder (MAE) framework [58], which was initially used to pretrain SAM’s original image encoder. SAMI follows an Encoder-Decoder pipeline: the encoder generates latent feature representations from unmasked patches, while the decoder reconstructs the missing embeddings of masked patches. This process is supervised by reconstruction loss, comparing the embeddings produced by SAM’s ViT-H encoder with those generated by the SAMI pipeline. After pretraining, the lightweight encoder is extracted from the SAMI pipeline and integrated with the rest of SAM’s components, forming EfficientSAM. The final step involves fine-tuning the entire model on the SA-1B dataset for further alignment and refinement. SAMI is a general pretraining approach that can be applied to train any backbone for SAM variants. The overall structure of SAMI and EfficientSAM is illustrated in Fig. 6.
@@
Concurrent with EfficientSAM, a new variant named RMP-SAM [150] is designed to achieve real-time, all-purpose segmentation, including panoptic segmentation (PS), video instance segmentation (VIS), and interactive segmentation (equivalent to the SegAny task). RMP-SAM retains SAM’s basic Encoder-Decoder architecture but incorporates more efficient components to enhance performance. For the encoder, RMP-SAM combines Feature Pyramid Networks (FPN) [173] with deformable convolutions [174] to extract features from both images and videos, while using a prompt encoder to embed visual prompts. In the decoder, RMP-SAM employs a three-stage pipeline utilizing novel pooling-based dynamic convolutions to refine the mask tokens. The tokens generated in each stage, along with the feature maps from the encoder, serve as inputs. These inputs are first processed by dynamic convolutions [175], followed by refinement using Multi-Head Self Attention (MHSA) and a Feed Forward Network (FFN). After the decoder, two additional prompt adapters are introduced to enhance interaction between the visual prompts and segmentation tokens. The final masks are generated by computing the inner product between the updated tokens and updated prompts. The architecture of RMP-SAM is illustrated in Fig. 7.
Figure 7: The architecture of RMP-SAM[150]. The encoder embeds images and videos to feature maps and encodes visual prompts as queries (tokens). The multi-stage decoder is to refine queries by interacting with feature maps. In the end, the feature maps and refined queries are fed into adapters to generate masks and classes.
Recently, Ravi et al. introduced the Segment Anything Model 2 (SAM 2) [34], an extension of the original SAM, with objective of delivering high-quality, real-time promptable segmentation across both images and videos. In image segmentation tasks, SAM 2 is reported to achieve higher accuracy and a 6× improvement in efficiency compared to the original SAM. This significant advancement is largely attributed to its efficient image encoder, Hiera [59], a hierarchical ViT that has been simplified from MViTv2 [176] by removing redundant components and utilizing the MAE framework for training. Hiera is a streamlined, purely transformer-based architecture that runs faster and delivers better accuracy than traditional ViTs in both image and video tasks.